foxey3:

Solve x3 = 64 over 27. ±8 over 3 8 over 3 ±4 over 3 4 over 3

glizxzy:

i got an answer but...

ThatGayGirl:

You are not supposed to give straight answers.

ThatGayGirl:

4/3 is correct however, would you like to know how to get that answer?

glizxzy:

its wrong any way

ThatGayGirl:

No, its correct.

foxey3:

ψ(._. )>

glizxzy:

i thought you werent supposed to give the answer

glizxzy:

so how did you find out the answer cause i can put my whole equation up

ThatGayGirl:

So can you tell me how to cube root the answer?

glizxzy:

x³=64/27 Cube root the answer ³√x³=³√64/27 x=4/3. Check: x=4/3 x³=(4/3)³ x³=64/27. As a result, the correct answer is 4/3 or 4 over 3. If you wonder why don't I choose the answer as ±4 over 3 because of this ±4/3 Choose positive first (4/3)²=64/27

glizxzy:

so

ThatGayGirl:

I am going to assume that this was copy and pasted.

glizxzy:

why because you cant understand it

glizxzy:

yw foxey